OpenStudy (anonymous):

what's the answer to this question? 2a3/2-3a5/2?

OpenStudy (mandre):

I presume you want to simplify it. There isn't much you can do with it. You could use a common denominator to make it one term.

OpenStudy (anonymous):

i didn't know how to do it,. can u help me?

OpenStudy (mandre):

\[\frac{ 2a ^{3} }{ 2 }-\frac{ 3a ^{5} }{2 }\] What is the common denominator

OpenStudy (anonymous):

2

OpenStudy (anonymous):

i didn't know how to factor it

OpenStudy (mandre):

No need to factor it. Now combine the 2 where 2 is the denominator and the numerator will be a combo of the 2 numerators.

OpenStudy (anonymous):

im going to multiply the denominators? and then as well as the numerators?

OpenStudy (mandre):

No \[\frac{ x }{ 2 }-\frac{ y }{ 2 }=\frac{x-y}{2}\] See what I did there?

OpenStudy (anonymous):

ahh,.. im going to subtract the numerator? then how about the 2 numbers there,. having a variable? what should i do with that?

OpenStudy (mandre):

You can't do anything with them. They stay the same. That will be the simplest form.

OpenStudy (anonymous):

ahh,. so the answer in the solution 2a3/2-3a5/2 is = 1a3-5/2?

OpenStudy (mandre):

No no no Your 2 terms stay the same x = 2a3 y = 3a5 Your answer is (x - y) / 2 2a3 stays the same and 3a5 stays the same.

OpenStudy (anonymous):

hha,.! tnx,.. but? can u answer that for me.? i really didn't understand,. im confused,..

OpenStudy (anonymous):

plzzzzz,.?? im so weak in math,.. i need your help,..

OpenStudy (anonymous):

\[\frac{ 2a ^{2} }{ 2 } - \frac{ 3a ^{5} }{2 }\]

OpenStudy (anonymous):

\[\frac{ 2a ^{2}-3a ^{5} }{ 2 }\]

OpenStudy (anonymous):

\[a^{2}- a^{5} = a^3\]

OpenStudy (anonymous):

\[\frac{ 2-3a^3 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 3a^3 }{ 1 }\]

OpenStudy (anonymous):

hopefully this helps